the conditions for the parabola y2 =4(x-d) & y2=4ax have a common normal other than x axis (a>0,c>0) is a>c>0)?
 

Grade:12

1 Answers

Riddhish Bhalodia
8 years ago
we have many equations
let the common normal intersect P1 at (x1,y1) and P2 at (x2,y2)

1) we have
y_1^2=4(x_1-d)
2) we have
y_2^2=4ax_2
3) By equating the nomal slope at points, and the slope of the line
we get
-y_2/2ax_2 = -y_1/2x_1 = (x_2-x_1)/(y_2-y_1)

eliminate any three variables and we get a quadratic in the remainig variable, it’s D>=0
that will give you the condition